OpenStudy (anonymous):

What do x and y equal? y=8-x 7=2-y

OpenStudy (texaschic101):

y = 8 - x now sub 8 - x in for y in the other equation 7 = 2 - y 7 = 2 - (8 - x) 7 = 2 - 8 + x 7 = -6 + x 7 + 6 = x 13 = x now sub 13 in for x in either of the equations to find y y = 8 - x y = 8 - 13 y = - 5 solution is : (13,-5)

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

do you have any questions ?

OpenStudy (anonymous):

No I think I got it

OpenStudy (texaschic101):

cool...glad I could help :)